OpenStudy (anonymous):

I need a easier understanding of subtracting integers my teacher gave me this note but I'm still having trouble. This is the note Same sign sum and keep different signs difference keep the sign of the larger number,then you'll be exact

OpenStudy (anonymous):

same sign add and keep the sign, different sign subtract and keep the sign of the bigger one. for example -9 - 8 = -9 + (-8) = you add and keep the sign = -17. For the other case -9 + 3 = you subtract and keep the sign of the greater one = -6

OpenStudy (anonymous):

Thanks @ dss08a

OpenStudy (anonymous):

no problem